# LVBench: An Extreme Long Video Understanding Benchmark

### v2

@@ 4 Experiments
4.1 Settings
We evaluated the performance of nine models that support multi images or short video input: TimeChat (Ren et al., 2023), Video-ChatGPT (Maaz et al., 2023), PLLaVA (Xu et al., 2024), LLaVA-Onevision (Li et al., 2024), CogVLM2-Video (Hong et al., 2024), LLaVA-NeXT (Zhang et al., 2024), GPT-4o (OpenAI, 2024), GLM4V-Plus (Hong et al., 2024) and InternVL2-40B (Chen et al., 2024b). To adapt these models for long video inputs, we sample a fixed number of frames from the original video, such as 32 or 96 frames, to maintain consistency with the model’s training sequence length. Additionally, we assessed six models that natively support long videos: LLaMA-VID (Li et al., 2023d), MovieChat (Song et al., 2023), LWM (Liu et al., 2024a), Gemini 1.5 Pro (Reid et al., 2024), Kangaroo (Liu et al., 2024b) and Qwen2-VL-72B (Wang et al., 2024). We processed the videos at a rate of 1 frame per second and fed them into the models, only performing downsampling when the video’s length exceeded the model’s maximum processing capability. It is worth noting that although Gemini 1.5 Pro can handle videos up to 10 hours long, its publicly available interface is limited to processing videos of up to 1 hour in length. For each question, we provided the following prompt as input to the models:
After obtaining the model responses, we first attempted to extract the answers using regular expression matching. For questions where the matching process was unsuccessful, we employed a GLM-4 model to extract the answers from the responses.
4.2 Evaluation Results
4.2.1 Performance across Core Capabilities
Overall, Qwen2-VL-72B achieved the best performance, outperforming other models in multiple tasks such as entity understanding (EU), temporal grounding (TG), reasoning (Rea) and summarization (Sum). Interestingly, some models that do not natively support long videos still managed to achieve competitive results compared to native long video support models. In terms of the overall score, InternVL2-40B ranked second only to Qwen2-VL-72B, and GLM4V-Plus ranked third, tying with the native long video support model Kangaroo.
However, the results of three widely used long video models, LLaMA-VID, MovieChat, and LWM, were nearly equivalent to random selection, highlighting the significant challenges that current models face when processing extremely long videos. This suggests that despite the ability to input long videos through model structure optimization, the performance and effectiveness of these models have not substantially improved.
Table 2: LVBench evaluation results regarding each core long video understanding capability. The highest score are highlighted with green, and the second highest are highlighted with purple. All the numbers are presented in % and the full score is 100%.
Model LLM ER EU KIR TG Rea Sum Overall
Non-Native Long Video Support Models
TimeChat (Ren et al., 2023) LLaMA2-7B 21.9 21.7 25.9 22.7 25.0 24.1 22.3
Video-ChatGPT (Maaz et al., 2023) Vicuna-1.5-13B 22.9 22.6 22.7 25.5 23.4 24.1 23.1
PLLaVA (Xu et al., 2024) Yi-34B 25.0 24.9 26.2 21.4 30.0 25.9 26.1
LLaVA-OneVision (Li et al., 2024) LLaMA3-70B 25.0 26.9 29.2 30.9 25.4 31.0 26.9
CogVLM2-Video (Hong et al., 2024) LLaMA3-8B 28.3 26.9 31.0 25.1 25.5 38.9 28.1
LLaVA-NeXT (Zhang et al., 2024) Yi-34B 30.1 31.2 34.1 31.4 35.0 27.6 32.2
GPT-4o (OpenAI, 2024) GPT-4 35.9 30.8 35.5 28.3 33.5 34.5 34.7
GLM4V-Plus (Hong et al., 2024) GLM-4 39.9 35.8 34.8 37.7 40.0 32.8 38.3
InternVL2-40B (Chen et al., 2024b) Nous-Hermes-2-Yi-34B 37.4 39.7 43.4 31.4 42.5 41.4 39.8
Native Long Video Support Models
MovieChat (Song et al., 2023) Vicuna-7B 21.3 23.1 25.9 22.3 24.0 17.2 22.5
LLaMA-VID (Li et al., 2023d) Vicuna-13B 25.4 21.7 23.4 26.4 26.5 17.2 23.9
LWM (Liu et al., 2024a) LLaMA2-7B 24.7 24.8 26.5 28.6 30.5 22.4 25.5
Gemini 1.5 Pro (Reid et al., 2024) Gemini 1.5 Pro 32.1 30.9 39.3 31.8 27.0 32.8 33.1
Kangaroo (Liu et al., 2024b) LLaMA3-8B 38.6 37.9 29.6 35.0 41.3 36.2 38.3
Qwen2-VL-72B (Wang et al., 2024) Qwen2-72B 38.0 41.1 38.3 41.4 46.5 46.6 41.3
4.2.2 Answer Distribution
Figure 4: Distribution of answers generated by different models.
To understand why some native long video support models perform poorly on LVBench, we evaluated the distribution of answers generated by different models on LVBench and observed that existing long video understanding models struggle with precisely following instructions. For example, despite explicitly constraining the output in the prompt to be one of four provided answer choices, Gemini 1.5 Pro generated responses outside of the specified options 20.9% of the time, such as "None of the above options are correct" or "I cannot answer this question". This occurred even though manual validation confirmed that the questions were indeed answerable from the given choices. MovieChat and LWM exhibited a strong bias towards selecting option A, regardless of the question. In contrast, InternVL2-40B demonstrated the strongest instruction-following capability, never generating responses outside the constrained options and producing a nearly uniform distribution over the answer choices.
We hypothesize that this discrepancy arises from the relatively higher quality and diversity of image-based instruction fine-tuning data compared to video instruction data. As InternVL2-40B ingests fewer image inputs, it can more readily generalize the learned capability of precisely following instructions from the image modality to video.
4.3 Performance across Video Categories
We conducted a comprehensive evaluation across various video categories. We selected two state-of-the-art models, InternVL2-40B and Qwen2-VL-72B, for testing and compared their results with human performance.
As shown in Table 3, humans achieved very high accuracy across all video categories, with an average of 94.4%. In contrast, the overall performance of InternVL2-40B and Qwen2-VL-72B was relatively lower, at only 39.5% and 41.3%, respectively. This indicates that there is still a significant performance gap between current multimodal models and humans in video understanding tasks, suggesting considerable room for improvement in understanding and analyzing long video content.
@@ 4.3 Performance across Video Categories
InternVL2-40B 43.5 45.2 38.9 41.6 32.8 36.4 39.5
Qwen2-VL-72B 43.0 42.6 40.8 41.0 42.0 38.9 41.3
Further analysis of the results for each category revealed that InternVL2-40B performed best on documentary videos, reaching an accuracy of 45.2%, while performing worst on TV shows, with only 32.8%. Qwen2-VL-72B, on the other hand, achieved the highest accuracy of 43.0% on sports videos and the lowest performance of 38.9% on cartoon.
4.4 Impact of LLM Filtering Method
Table 4: Ablation study on LLM filtering method.
Model w/ LLM w/o LLM
LWM 25.5 32.7
LLaVA-NeXT 32.2 48.9
Table 4 demonstrates the effectiveness of using large language models to filter question-answer pairs. Despite instructing the annotators to watch the entire video before labeling, a significant proportion of questions could still be answered correctly by inferring from the matching degree between the question and options, as well as the differences among the options. The experimental results show that after applying the LLM filter, the score of LWM decreased from 32.7% to 25.5%, while the score of LLaVA-NeXT, which employs a more powerful language model, dropped from 48.9% to 32.2%, with a decline of 16.7 percentage points. This indicates that stronger language models have a higher probability of inferring the correct answer solely from the natural language context, highlighting the importance of the data filtering step in the process.
4.5 Impact of Video and Clue Length
We investigated the impact of different video durations and clue durations on the experimental results. As shown in Figure 5, the performance of GLM4V-Plus, InternVL2-40B, and Qwen2-VL-72B remains relatively stable across various video length intervals, demonstrating overall strong performance.
Clue durations refers to the time span of video content needed to answer a specific question. Figure 5 illustrates that most models perform well when the clue duration is between 0-10 seconds or greater than 60 seconds. This may be attributed to the fact that questions with clues longer than 60 seconds tend to focus more on analyzing and summarizing the relationships between multiple events. These models are equipped with stronger language modeling capabilities, giving them an advantage in tackling such problems.
Figure 5: The impact of different video and clue durations.
@@ Conclusion.
In this paper, we introduced LVBench, a benchmark designed to advance long video understanding. LVBench comprises a diverse collection of lengthy videos and a meticulously annotated question-answer dataset, presenting a robust evaluation framework for assessing multimodal models on complex video understanding tasks. Our experiments revealed that while state-of-the-art models have made strides in short video understanding, their performance on long videos falls short of human-level accuracy. By providing a challenging benchmark, we hope to stimulate the development of advanced models capable of tackling the complexities of extended video comprehension.
Limitations.
@@
A limitation of our benchmark is the exclusion of audio data. While audio can provide valuable context, we did not include it because most current models lack effective audio processing capabilities. Future work will aim to incorporate audio information to enhance the evaluation framework for multimodal understanding.

### v3

@@ 4 Experiments
Seed1.5-VL 63.3 67.5 60.3 66.9 60.4 65.0 64.0
Gemini-2.5-Pro 71.3 54.3 72.1 68.5 69.2 66.1 67.4
4.1 Settings
We evaluated the performance of 13 models that support multi images or short video input: TimeChat [33], Video-ChatGPT [23], PLLaVA [43], LLaVA-OneVision [16], CogVLM2-Video [12], LLaVA-NeXT [50], InternVL2-40B [6], mPLUG-Owl3 [44], GLM-4.1V-Thinking [13], VideoLLaMA3-7B [47], GLM4V-Plus [12], GPT-4o [26] and GPT-4.1 [27]. To adapt these models for long video inputs, we sample a fixed number of frames from the original video, such as 32 or 96 frames, to maintain consistency with the model’s training sequence length. Additionally, we assessed 13 models that natively support long videos: LLaMA-VID [19], MovieChat [34], LWM [20], Gemini 1.5 Pro [32], Kangaroo [21], Qwen2-VL-72B [37], Qwen2.5-VL-72B [4], Gemini-2.0-Flash [7], AdaReTaKe [39], Gemini-2.5-Flash [7], MR.Video [28], Seed1.5-VL [10] and Gemini-2.5-Pro [7]. We processed the videos at a rate of 1 frame per second and fed them into the models, only performing downsampling when the video’s length exceeded the model’s maximum processing capability. It is worth noting that although Gemini 1.5 Pro can handle videos up to 10 hours long, its publicly available interface is limited to processing videos of up to 1 hour in length. For each question, we provided the following prompt as input to the models:
After obtaining the model responses, we first attempted to extract the answers using regular expression matching. For questions where the matching process was unsuccessful, we employed a LLM to extract the answers from the responses.
4.2 Performance across Core Capabilities
Overall, Gemini-2.5-Pro demonstrated state-of-the-art (SOTA) performance, achieving a top score of 67.4 by outperforming all other models in 5 out of 6 tasks (EU, KIR, TG, Rea, and Sum). Seed1.5-VL-Thinking secured the second position with a score of 64.6. Our analysis reveals two key findings. First, a significant performance gap persists between proprietary and open-source models. The top-performing open-source models, VideoLLaMA3-7B (non-native long-video) and Qwen2.5-VL-72B (native long-video), scored 45.3 and 44.0, respectively, lagging behind Gemini-2.5-Pro by over 20% in both individual tasks and the overall score. Second, we observed that some models without native long-video support still achieved competitive results against their native counterparts, though their performance was not comparable to SOTA models. Conversely, several models designed for native long-video processing, including MovieChat, LLaMA-VID, LWM, and Gemini-1.5-Pro, exhibited unexpectedly low scores.
4.3 Analysis of Failure Modes
To understand why some native long video support models perform poorly on LVBench, we evaluated the distribution of answers generated by different models on LVBench and observed that existing long video understanding models struggle with precisely following instructions. For example, despite explicitly constraining the output in the prompt to be one of four provided answer choices, Gemini-1.5-Pro generated responses outside of the specified options 20.9% of the time, such as ”None of the above options are correct” or ”I cannot answer this question”. This occurred even though manual validation confirmed that the questions were indeed answerable from the given choices. MovieChat and LWM exhibited a strong bias towards selecting option A, regardless of the question.
We hypothesize that although these models have the ability to process long videos, they may lack instruction data for such long videos, which leads to them being unable to effectively follow instructions, in turn affecting their performance.
4.4 Performance across Video Categories
We conducted a comprehensive evaluation across various video categories. We selected two leading models, Seed1.5-VL and Gemini-2.5-Pro, for detailed analysis and compared their results with human performance.
As shown in Table 3, humans achieve a very high accuracy of 94.4% on average, setting a strong benchmark across all categories. In contrast, the overall performance of Gemini-2.5-Pro and Seed1.5-VL was considerably lower, at 67.4% and 64.0%, respectively. This highlights a significant performance gap between current multimodal models and humans in long-video understanding, indicating substantial room for improvement.
Further analysis of the results by category reveals distinct model strengths. Gemini-2.5-Pro performed best on Event Record videos, reaching an accuracy of 72.1%, but struggled with Documentary videos, where its score dropped to 54.3%. Conversely, Seed1.5-VL demonstrated more consistent performance across categories, with all scores falling within a narrow range between 60.3% (Event Records) and 67.5% (Documentaries).
4.5 Ablation on Frame Density
To understand the relationship between temporal information density and model performance, we conduct an ablation study on the number of input frames. We evaluate three state-of-the-art models—Seed1.5-VL, Gemini-2.5-Pro, and Qwen2.5-VL-72B—with varying frame inputs: 0, 1, 4, 8, 50, and a dense sampling of 1 frame per second (FPS). The results are presented in Figure  4.
As depicted in the figure, all models exhibit a clear trend of improved performance with an increased number of input frames. A modest performance gain is observed as the frame count increases from 1 to 8. This suggests that with sparse frames, models can grasp the general context or primary event of the video but may fail to capture the transient, crucial visual cues necessary for accurate reasoning. The most significant finding is the substantial performance leap from 50 frames to the 1 FPS setting. This large gap underscores the necessity of dense visual input for resolving the complex, long-range temporal dependencies inherent in our benchmark tasks.
Furthermore, the “0 Frames” condition serves as a critical control experiment to validate the integrity of our benchmark’s question-answer pairs. In this visuals-agnostic setting, models must attempt to answer based solely on linguistic cues from the question and options. The results show that all models perform near the level of random guessing. This outcome effectively demonstrates that success on our benchmark is contingent on visual understanding rather than exploitable language biases. It also validates the efficacy of our LLM-based data filtering protocol, which successfully purges questions that do not strictly require visual grounding.
Figure 4: Impact of Frame Count on Model Performance.
@@ 5 Discussion
Conclusion. In this paper, we introduced LVBench, a benchmark designed to advance long video understanding. LVBench comprises a diverse collection of lengthy videos and a meticulously annotated question-answer dataset, presenting a robust evaluation framework for assessing multimodal models on complex video understanding tasks. Our experiments revealed that while state-of-the-art models have made strides in short video understanding, their performance on long videos falls short of human-level accuracy. By providing a challenging benchmark, we hope to stimulate the development of advanced models capable of tackling the complexities of extended video comprehension.
@@
Limitations. A limitation of our benchmark is the exclusion of audio data. While audio can provide valuable context, we did not include it because most current models lack effective audio processing capabilities. Future work will aim to incorporate audio information to enhance the evaluation framework for multimodal understanding.
License. Our dataset is under the CC-BY-NC-SA-4.0 license. The dataset relies on YouTube as an external resource for the video content. We only provide download links rather than source videos. While there is no guarantee that the original YouTube videos will remain constant over time, we have taken measures to mitigate this issue by creating and storing copies of the videos. These copies can be provided to dataset consumers, subject to obtaining permission from the original video creators.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants 62425601 and 62495063, and by the New Cornerstone Science Foundation through the XPLORER PRIZE.
Appendix A Datasheet
A.1 Motivation
  • For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
    The dataset was created to evaluate and optimize long video understanding models. To the best of our knowledge, this is the first benchmark capable of assessing models’ performance on extra-long videos.
  • Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
    Anonymous for review
  • Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
    Anonymous for review
  • Any other comments?
    No.
A.2 Composition
  • What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
    The instances comprising the dataset are YouTube videos, which can be categorized into six different types: Sports, Documentary, Event Record, Lifestyle, TV Show, and Cartoon.
  • How many instances are there in total (of each type, if appropriate)?
    In total, there are 103 videos and 1,549 question-answer pairs in the dataset.
  • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
    The dataset contains a sample of instances from a larger set of YouTube videos and their corresponding question-answer pairs. The multi-stage filtering process was employed to select high-quality question-answer pairs for inclusion in the dataset.
  • What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
    Each instance in the dataset consists of a single video and its corresponding supplementary information. All the additional information is stored in the “video_info.meta.jsonl” file, where:
    • All question-answer pairs associated with the video are located under the “qa” key.
    • The video category is specified under the “type” key.
    • Details such as video duration, resolution, and other metadata are contained within the “video_info” key.
  • Is there a label or target associated with each instance? If so, please provide a description.
    Yes, each video instance in the dataset is associated with a set of corresponding question-answer pairs, which serve as the labels or targets for that particular instance.
  • Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
    No.
  • Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
    N/A.
  • Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
    No, there are no recommended data splits for this dataset. The entire dataset is intended to be used for evaluating and testing the performance of long video understanding models.
  • Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
    Yes, there is a small probability of bias introduced by the annotators during the labeling process.
  • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a dataset consumer? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
    The dataset relies on YouTube as an external resource for the video content. While there is no guarantee that the original YouTube videos will remain constant over time, we have taken measures to mitigate this issue by creating and storing copies of the videos. These copies can be provided to dataset consumers, subject to obtaining permission from the original video creators.
  • Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
    No.
  • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
    No.
  • Any other comments?
    No.
A.3 Collection Process
  • How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If the data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
    The data associated with each instance in the dataset was manually annotated by hired annotators.
  • What mechanisms or procedures were used to collect the data (e.g., hardware apparatus(es) or sensor(s), manual human curation, software program(s), software API(s))? How were these mechanisms or procedures validated?
    The data collection process involved the use of dedicated servers to download the YouTube videos, followed by the annotation of the videos using an annotation platform.
  • If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
    See the third question of the previous section on “Composition”.
  • Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
    The data annotation process was carried out by professional annotators, who were compensated with an average of 30 US dollars per video annotated.
  • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
    Anonymous for review
  • Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
    No.
  • Any other comments?
    No.
A.4 Preprocessing/cleaning/labeling
  • Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remaining questions in this section.
    Yes, preprocessing and cleaning of the data was performed. We manually inspected the content of each video to ensure all videos were high-quality, content-rich, and free of harmful information.
  • Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
    No.
  • Is the software that was used to preprocess/clean/label the data available? If so, please provide a link or other access point.
    No.
  • Any other comments?
    No.
A.5 Uses
  • Has the dataset been used for any tasks already? If so, please provide a description.
    Not beyond this paper.
  • Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
    Anonymous for review
  • What (other) tasks could the dataset be used for?
    The dataset could be used for various other long video understanding tasks such as summarization, captioning, question answering and multimodal understanding.
  • Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)? If so, please provide a description. Is there anything a dataset consumer could do to mitigate these risks or harms?
    No.
  • Are there tasks for which the dataset should not be used? If so, please provide a description.
    No.
  • Any other comments?
    No.
A.6 Distribution
  • Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
    Yes, the dataset is freely and publicly available and accessible.
  • How will the dataset be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
    Anonymous for review
  • When will the dataset be distributed?
    Anonymous for review
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
    Our dataset is under the CC-BY-NC-SA-4.0 license.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
    No.
  • Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
    No.
  • Any other comments?
    No.
A.7 Maintenance
  • Who will be supporting/hosting/maintaining the dataset?
    Anonymous for review
  • How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
    Anonymous for review
  • Is there an erratum? If so, please provide a link or other access point.
    No.
  • Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to dataset consumers (e.g., mailing list, GitHub)?
    Anonymous for review
  • If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
    Yes, there are applicable limits on the retention of data associated with individuals in the videos. If the original creators of the videos believe their content should be removed from the dataset, we will remove it upon request.
  • Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
    Anonymous for review
  • If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description.
    No, there is no mechanism for others to extend, augment, build on, or contribute to the dataset. This decision was made to ensure consistency and fairness in benchmarking and evaluation using the dataset.
  • Any other comments?
    No.
Figure 5: The proportion of core capability combinations.
Appendix B Distribution of Core Capability Combinations
In this section, we quantified the distribution of various combinations of core competencies within the dataset. As illustrated in the Figure 5, the six core competencies can be further combined to form 26 fine-grained question types. This flexible combinatorial approach guarantees the richness and diversity of the dataset, enabling a comprehensive evaluation of the model’s performance across multiple dimensions.